The perimeter of a rectangular piece of metal is 44 centimetres. The area is 112 square centimetres. What are the dimensions of the piece?

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

The perimeter of a rectangular piece of metal is 44 centimetres. The area is 112 square centimetres. What are the dimensions of the piece?
----------------------
P = 2L + 2W = 44
L + W = 22
---
L*W = 112
---
Hint: L & W are integers

Answer by ikleyn(52781)

.
This problem is for MENTAL solution.

Find two co-factors of 112 that sum up to 22.